in triangle abc, the measure of angle b is 11 degrees more than three times the measure of angle a. the measure of angle c is 49 degrees more than the measure of angle a. I have been stuck on this for too long, please help!
Answer by Alan3354(69443)

in triangle abc, the measure of angle b is 11 degrees more than three times the measure of angle a. the measure of angle c is 49 degrees more than the measure of angle a.
------------------
b = 3a + 11
c = a + 49
--------------
a + b + c = 180
a + (3a+11) + (a+49) = 180
Can you go from there?